OpenStudy (jake124):

find the quotient 0.06 divided by 7.8 A.130 B. 0.013 C. 1.3 D. 0.13 PLs help WIL MEDAL AND FAN !!!!

OpenStudy (hocuspucus292333393):

Its A.

OpenStudy (mathmale):

A preliminary step that may make this work a bit easier would be to eliminate the decimal points from the two given numbers. This is optional, not mandatory. Multiply 0.06 by 100, and then multiply 7.8 by 100 also. Your result? Since 780 is larger than 6, divide 6 into 780. Show your work and result.

OpenStudy (sshayer):

\[\frac{ .06 }{ 7.80 }=\frac{ 6 }{ 780 }=\frac{ 1 }{ 130 }=?\]

OpenStudy (sshayer):

from your options it seems question should be divide 7.8 by .06

OpenStudy (jake124):

thank you @ hocuspucus

OpenStudy (mathmale):

In summary, you could leave your result as 1/130, or you could divide 130 into 1. But the latter operation is not "nice;" you obtain a fraction that never stops.

OpenStudy (mathmale):

Not only is "HocusPucus292333393 Best Response Medals 0 "It's A," it's a "direct answer," with no involvement of the student in the solution of the problem. Kindly give out no more info than is necessary to enablet he student to find his or her own solution.

OpenStudy (mathmale):

@jake124: it'd be very wise not to accept answers from others. First of all, it violates the rules; second, you won't know whether the given answer is correct or not, nor how to find the correct answer yourself.